OpenStudy (anonymous):

Bea rolls a ball off the edge of a swimming pool and strikes the water 2.0 m away 0.30 sec. later. What was the initial speed of the ball?

OpenStudy (anonymous):

help @Akpellet_math

OpenStudy (anonymous):

help @QWERTYCraft

OpenStudy (anonymous):

help me @nincompoop

OpenStudy (anonymous):

help me @Darrius

OpenStudy (anonymous):

help me @thomaster

OpenStudy (anonymous):

Help me pls @wolfe8

OpenStudy (anonymous):

help me pls @hba

OpenStudy (anonymous):

help me @Akpellet_math @Albany_Goon @ZeHanz @eliassaab @OtonoGold

OpenStudy (anonymous):

help me @ememlove

OpenStudy (anonymous):

help me pls @Vinaire @Mashy @Mathbreaker @Chineseboy15 @CyberShadow @sisco0123 @StrikeEagle

OpenStudy (anonymous):

use equations of motion v=0 s=2m t=0.3 sec.

OpenStudy (anonymous):

Not enough information to conclude, we need to know the acceleration and inclination of the plane if it is so. Otherwise if youre talking about average speed then its 4 m/s

OpenStudy (anonymous):

so first use this v=u+at 0=u+a(0.3) a=-10u/3 now use ..v^2=u^2+2as 0=u^2+2(-10u/3)2 simplify and obtain u u=40/3 m/s u=13.33 m/s

OpenStudy (anonymous):

THANK YOU @Akpellet_math AND @Mathbreaker

OpenStudy (anonymous):

wait what? :O "now use ..v^2=u^2+2as 0=u^2+2(-10u/3)2" how is that? :O

OpenStudy (anonymous):

I dont think you can assume that final velocity is 0 akpellet

OpenStudy (anonymous):

The question is a lil absurd.. is it just falling off the edge? or rolling along an incline?

OpenStudy (anonymous):

@Mashy equations of motion ...and at last the velocity is becoming zero only .!

OpenStudy (anonymous):

@Akpellet_math .. no :P.. you just can't use equation and plug values like that .. besides when it hits the water, its velocity won't be zero.. think about it ;-)

OpenStudy (anonymous):

Just because it strikes the water doesnt mean it becomes 0 :P

OpenStudy (anonymous):

@Socallmesheckyy hey you got some diagram or something? :P

OpenStudy (anonymous):

Its impossible to get conclusive solution without more variables given.

OpenStudy (anonymous):

needs to be assumed that ..! otherwise then there's no way to the problem .! i know ..the solution is only mathematical ..!

OpenStudy (anonymous):

so do u have any proceeding to the problem .. i know ..my answer seems ..practically absurd .!

OpenStudy (anonymous):

it can be calculate @mathbreaker if its just falling off.. like this